### v4

# Harnessing natural and mechanical airflows for surface-based atmospheric pollutant removal Centre for Climate Repair, Cambridge, UK

Harnessing natural and mechanical airflows for surface-based atmospheric pollutant removal
Centre for Climate Repair, Cambridge, UK
Department of Mechanical Engineering, Stanford University, Stanford, USA
May 22, 2025)
Removal strategies for atmospheric pollutants are increasingly being considered to mitigate global warming and improve public health. However, surface-based removal techniques, such as sorption, catalysis and filtration, are often limited by pollutant transport and removal rate constraints. We evaluate the atmospheric pollutant transport to surfaces and assess the potential of surface-based removal technologies for applications in airflow through cities, HVAC systems and over vehicles. If these removal technologies are applied to their surfaces, cities, solar farms, HVAC systems and filters can achieve atmospheric pollutant removal rates that exceed 1 GtCO2e annually (20-year GWP). Cities have the highest atmospheric pollutant removal potential, with estimates averaging 30 GtCO2, 0.06 GtCH4, 0.0001 GtPM2.5, 0.007 GtNO annually. HVAC filters can achieve atmospheric pollutant removal costs as low as $300 per tCO2e removed when sorption or catalyst technologies are incorporated into their fibre sheets, outperforming the $2000 per tCO2e removal cost when these technologies are applied to city surfaces. This estimate is based on the literature values for these technologies’ costs per square meter. However, our calculations indicate that optimising catalyst properties and surface coverage could lower the cost estimates to below $100 per tCO2e across these applications. These findings demonstrate that integrating surface-based pollutant removal technologies into infrastructure may offer a scalable pathway to advance climate and health objectives.
Greenhouse gases (GHGs) such as carbon dioxide (CO2), methane (CH4) and nitrous oxide (N2O) continue to rise at significant rates, presenting sustainability and health challenges by accelerating climate change [22, 21]. In 2023, CO2 levels surpassed 420 ppm, driven by fossil fuel use and deforestation, contributing to biodiversity loss and environmental degradation [35, 22]. The situation is compounded by CH4 and N2O, which, although emitted in smaller quantities, have global warming potentials (GWPs) 84 and 273 times that of CO2 over a 20-year timescale. These potent GHGs are emitted by anthropogenic sources such as agriculture and fossil fuel processing [3, 56, 12, 38]. However, unlike CO2, CH4 and N2O have biogenic sources that are difficult to control without disrupting agriculture and ecosystems. In addition to GHGs, atmospheric pollutants such as particulate matter (PM), sulphur dioxide (SO2) and nitrogen oxides (NO) pose environmental and medical risks [27, 30]. PM, particularly particles smaller than 2.5 m (PM2.5), is a leading cause of respiratory and cardiovascular diseases, causing millions of premature deaths annually [47, 30, 33]. PM is emitted from anthropogenic sources, such as fossil fuel processing, and natural sources, like wildfires; SO2 and NO have similar sources, but contribute to acid rain and ground-level ozone [7, 2, 5].
Given the diffuse nature of anthropogenic, biogenic and natural sources, coupled with the insufficient progress in emissions reductions, there is growing interest in atmospheric pollutant removal [35, 22, 12, 38, 30, 33, 7, 2, 5]. Atmospheric pollutant removal can generally be divided into two categories: surface-based approaches [39, 25, 26, 16, 23, 31, 57, 55], which rely on the interaction between pollutants and surfaces, and airborne methods, which use radicals or surface chemistry to break down pollutants in the air [54, 36]. While airborne methods show promise, this study focuses on surface-based removal methods, given that these approaches are more widely explored and developed. Despite the interest in surface-based atmospheric pollutant removal [39, 25, 26, 16, 23, 31, 57, 55], the scale of potential removal remains unexplored. This study aims to assess the potential of various surface-based strategies by quantifying the scales at which pollutants can be removed from the atmosphere. Nonetheless, the methods developed here apply to other pollutants, offering a framework for quantifying the removal potential across contexts.
Direct air capture (DAC) technologies, often sorbent (e.g., amines) or solvent-based (e.g., potassium hydroxide), can capture CO2 through chemical adsorption and absorption processes [39, 8]. For DAC, the fraction of pollutant molecules removed upon contact with the surface, known as the removal efficiency, ranges from to [46]. However, this process requires substantial energy and involves high operational costs [29, 28]. Membrane-based systems, which separate CO2 through selectively permeable barriers, offer greater energy efficiency than DAC, but face challenges related to initial costs and scalability [25]. Hence, while both methods show promise, their widespread application has been limited. In contrast, leveraging pre-existing natural and mechanical airflows, where the associated energy demands are already accounted for, may enable scalable and cost-effective surface-based methods for atmospheric pollutant removal.
Catalytic oxidation can convert CH4 into CO2 and water (H2O), reducing its greenhouse impact; however, performance optimisation across environments remains challenging [26]. Reported CH4 removal efficiencies for thermal, photo and electrocatalysts vary widely, from to , reflecting differences across catalysts and experimental conditions [50, 32]. Recent cost analyses suggest that using photocatalytic oxidation for atmospheric CH4 removal is prohibitively expensive due to the high flow rates required for effective removal [36, 32, 18]. Alternatively, atmospheric radicals can break down CH4 but at slow, condition-dependent rates [54, 36] and methanotrophic bacteria can remove CH4, albeit under controlled conditions [16]. This study examines catalytic oxidation as a strategy for atmospheric pollutant removal, exploring applications with varying surface fluxes and areas for scalable implementation.
Similar to CH4, N2O can be removed via catalytic reduction, which converts N2O into nitrogen (N2) and oxygen (O2) [36]. Another surface-based removal process is biological denitrification, where bacteria reduce N2O to N2 under anaerobic conditions [49].
High-efficiency particulate air (HEPA) filters are effective at capturing PM [23], alongside other methods such as electrostatic precipitators and adsorption materials such as activated carbon [31]. These technologies offer removal efficiencies of to [42], but require periodic regeneration or replacement due to saturation [24]. PM is also removed from the atmosphere through dry (e.g., gravitational settling, surface absorption) and wet deposition (e.g., scavenging by rain), with removal rates varying by region and environmental conditions [45, 53, 15]. While these mechanisms for PM removal are well-established, our study will estimate how much PM2.5 is removed from the atmosphere by filtration technologies and dry deposition via surface absorption, comparing these with the potential removal rates of other pollutants.
Similar to CH4 and N2O, SO2 and NO can be removed using catalysis, which converts these pollutants into sulfuric acid (H2SO4) and N2 [57]. Another removal method is scrubbing, where SO2 and NO are removed from gas streams by passing them through chemically reactive liquid solutions [55].
In summary, although many surface-based pollutant removal methods exist (or are being proposed), they are often inefficient, energy-intensive and/or costly. Furthermore, the literature lacks an assessment of the potential scale of reductions achievable with surface-based atmospheric pollutant removal, despite the need for scalable solutions.
In this study, we examine approaches that could utilise the substantial volumes of air transported in urban areas, heating, ventilation and air conditioning (HVAC) systems, and transportation networks, where pollutants interact with surfaces to enable their removal. Rather than focusing on the performance of specific surface-based technologies that remove specific pollutants, this study generally evaluates the scalability of atmospheric pollutant removal via surface-based interactions through transport limitations. We use a combination of theoretical models, empirical relations, global measurements and industry standards to estimate the atmospheric pollutant flow rate to the surfaces of various natural and mechanical systems. By leveraging removal efficiencies measured in the literature [46, 50, 32, 42], we assess the projected removal rates and technology costs for sorption, catalysis and filtration, for a range of environmental and flow conditions. We compare the potential atmospheric pollutant removal rates with intergovernmental targets and existing removal mechanisms, evaluating the scalability and cost-effectiveness of these surface-based technologies.
Results
Natural (e.g., cities, solar farms), internal (e.g., HVAC, combustion, DAC systems) and external environments (e.g., aeroplanes, trains, automobiles) offer significant opportunities for atmospheric pollutant removal. As shown in Fig. 1a and detailed in Supplementary Sec. 3, these environments consist of substantial airflow volumes that drive atmospheric pollutant fluxes to surfaces. Average streamwise flow rates are approximately m3/s through natural environments, m3/s in internal systems and m3/s around external systems. Theoretical predictions, empirical data, global measurements and industry standards provide a range of estimates in Fig. 1a, spanning from to m3/s for cities, to m3/s for HVAC systems and to m3/s for automobiles. Key parameters, including length, surface area, velocity and the number of environments are summarised in Fig. 1b–g, Supplementary Tabs. 2–4 and Secs. 2.1–2.3. The parameter distributions not only capture uncertainty but reveal that surface area and airflow velocity are the dominant factors influencing pollutant removal efficiency across applications.
Surface fluxes of atmospheric pollutants
Fig. 2a shows normalised atmospheric pollutant fluxes to the surfaces of natural environments (columns 1–2) in cities and solar farms. GHG fluxes exceed other pollutant fluxes due to higher diffusivities and atmospheric concentrations (Fig. 2c). Atmospheric pollutant fluxes to surfaces are also influenced by length scales, with cities exhibiting average surface fluxes of mol/m2s for CO2, while solar farms achieve average surface fluxes of mol/m2s for CO2. Smaller length scales reduce the boundary layer thickness, which increases the turbulence intensity and improves the atmospheric pollutant transfer to surfaces. As a result, in these natural environments, atmospheric pollutant fluxes to surfaces are averaged across scales. Using empirical formulae detailed in the Methods, average atmospheric pollutant fluxes at the repeating unit scale (e.g., to the surface of individual buildings, m-1) are larger than those at the combined unit scale (e.g., to the surface of entire cities, m-1) due to steeper concentration gradients (Figs. 2a, d, e).
Figs. 2a–b highlight the enhanced atmospheric pollutant transport to the surfaces of internal airflows (columns 3–5) in HVAC, combustion, DAC systems (Fig. 2a) and HVAC filters (Fig. 2b). Comparing HVAC, combustion and DAC systems, which show average surface fluxes of , and mol/m2s for CO2 in Fig. 2c, demonstrates that higher airflow velocities enhance atmospheric pollutant transport. The estimated value for DAC is comparable to mol/m2s estimated in McQueen et al. [28]. At the duct wall, the rate at which pollutants move perpendicular to the surface (normal flux) depends on diffusion and is proportional to (e.g., an average of mol/m2s for CO2, Fig. 2c, f), where is the diffusivity of the pollutant, is its concentration and is the height of the channel. In contrast, for HVAC filters, the pollutant flux through the porous fibre sheet is proportional to (e.g., an average of mol/m2s for CO2, Fig. 2c, g), where is the airflow velocity through the fibre sheet. Since and are typically small relative to and , is generally larger than , highlighting the benefits of utilising advective over diffusive transport to improve atmospheric pollutant fluxes to surfaces.
Lastly, we evaluate the normalised atmospheric pollutant fluxes to the surfaces of external airflows over aeroplanes, trains and automobiles (columns 6–8, Fig. 2a). Automobiles demonstrate higher pollutant fluxes to surfaces than trains and aeroplanes (e.g., an average of , , mol/m2s for CO2, Fig. 2c), due to smaller characteristic lengths. However, aeroplanes generate more turbulent transport due to their higher velocities, resulting in a pollutant flux that is approximately 25% of the automobile flux across the entire surface.
Surface flow rates of atmospheric pollutants
Fig. 3 quantifies the flow rates of atmospheric CO2, CH4, PM2.5 and NO to all global surfaces. These flow rates are evaluated by multiplying the fluxes from Fig. 2 by the relevant surface area for each application ( in Fig. 2d–g) and the number of each application. Natural and internal airflows (columns 1, 3, 4) with extensive surface areas yield the highest atmospheric pollutant flow rates to their surfaces, e.g., an average of 30 GtCO2/y for cities, 10 GtCO2/y for HVAC systems and 4 GtCO2/y for HVAC filters. Conversely, external airflows and small-surface-area internal systems (columns 5–8) show lower atmospheric pollutant flow rates to their surface, e.g., an average of 0.1 GtCO2/y for aeroplanes, 0.7 GtCO2/y for trains, GtCO2/y for combustion systems and GtCO2/y for DAC systems, despite external airflows exhibiting higher atmospheric pollutant fluxes to their surfaces in Fig. 2a–c.
Integrating surface-based pollutant removal technologies into global infrastructure is likely more readily achieved by incorporating technologies into newly-produced surfaces. Comparing atmospheric pollutant flow rates to total existing surfaces (Fig. 3) with annually produced surfaces (Supplementary Fig. 2), a decrease is observed across applications, with the reduction varying between applications. For example, solar farms and cities have average flow rates of 30 GtCO2/y and 0.08 GtCO2/y to their total surfaces. However, when accounting for annually produced surfaces, the average flow rate to new surfaces in cities (0.2 GtCO2/y) is more significantly reduced than for solar farms (0.03 GtCO2/y). This is largely driven by the global growth of solar farms, expanding at approximately 20% annually, compared with the 4% annual growth of cities.
Removal rates of atmospheric pollutants
Having established the atmospheric pollutant transport rates to the surfaces of various natural and mechanical environments, we assume ideal removal efficiencies and compare the potential atmospheric pollutant removal rates with the 1 GtCO2e/y target (20–100-year GWP) [20]. This target appears feasible for cities, solar farms, HVAC systems, filters, aeroplanes, trains and automobiles, based on the potential atmospheric pollutant flow rates to existing surfaces (Fig. 3). When focusing only on potential atmospheric pollutant flow rates to annually produced surfaces (Supplementary Figs. 2), feasible applications narrow to cities, solar farms, HVAC systems and filters.
| Carbon dioxide (CO2) | |||||
| Environment | Rem. rate (t/y) | Surf. area (m2/y) | Tech. cost ($/y) | Rem. cost ($/t) | Rem. cost ($/tCO2e) |
| Cities | |||||
| Solar farms | |||||
| HVAC systems | |||||
| HVAC filters | |||||
| Methane (CH4) | |||||
| Environment | Rem. rate (t/y) | Surf. area (m2/y) | Tech. cost ($/y) | Rem. cost ($/t) | Rem. cost ($/tCO2e) |
| Cities | |||||
| Solar farms | |||||
| HVAC systems | |||||
| HVAC filters | |||||
However, the removal efficiency is often suboptimal, falling short of the transport flux to surfaces. For example, DAC systems exhibit removal efficiencies around [46], such that the average pollutant removal potential is 1 GtCO2/y for HVAC systems. With 10 DAC systems removing, on average, around tCO2/y each [28], atmospheric removal rates are currently limited to approximately GtCO2/y (Fig. 3a). CH4 and N2O present additional challenges due to their low atmospheric concentrations but can also achieve removal efficiencies around [50]. However, most efficiency values are reported at elevated temperatures, with room temperature conversion rates still being an area of active research. When multiplied by the removal efficiency, cities, solar farms and HVAC systems have maximum pollutant removal potentials (20, 0.5, 2 GtCH4/y respectively), which are greater than atmospheric oxidation (0.3 GtCH4/y), providing a pathway that could surpass current emissions [9, 49]. PM2.5 is captured by HVAC filters which achieve removal efficiencies of around [42], such that the average pollutant removal potential is 0.01 GtPM2.5/y for HVAC filters. Atmospheric removal rates approach GtPM2.5/y (Fig. 3c), with a mean estimate of 100 million filters processing 100 m3/h at 99% efficiency [24].
Cost analyses
Finally, we estimate the potential cost of scaling CO2-sorption and CH4-catalyst technologies, highlighting their capacity for atmospheric pollutant removal across natural and mechanical airflows. These surface-based technologies are selected as promising approaches for GHG removal, with estimated costs ranging from $17 to $73 per m2 for CO2 sorbents and $14 to $60 per m2 for catalysts targeting CH4 based on [39, 29, 36, 18]. By contrast, atmospheric PM2.5, as well as the oxidised forms of NO, are captured when they interact with surfaces through deposition. As shown in Fig. 3c–d, dry deposition rates of GtPM2.5/y and GtNO/y align with atmospheric pollutant flow rate estimates to urban surfaces [53, 10], contributing around 1–10% of global deposition.
To evaluate the potential cost-efficiency of each application, we compare relative to atmospheric pollutant removal costs of $100 per tCO2e removed (20-year GWP), consistent with recommendations for GHG removal technologies [20]. We assume a removal efficiency of [46, 50, 32, 24], reflecting a realistic and optimistic average across surface-based removal technologies and pollutants. By dividing the estimated annual surface-based technology cost by the potential atmospheric pollutant removal rate, we derive the cost per tonne of CO2e removed for each application in Tab. 1. Estimates across natural and internal environments exceed the $100 per tCO2e removed target. However, HVAC filters achieve potential costs as low as $300 per tCO2 removed for sorption and $600 per tCO2e removed for catalyst technologies. This lower cost for HVAC filters is attributed to the high atmospheric pollutant flux to their surface (e.g., an average of mol/m2s for CO2, Fig. 2c), which reduces the influence of surface area on the potential removal cost.
We explore the potential of different catalysts for atmospheric methane removal by varying material properties and surface coverage. Palladium (Pd) and magnesium (Mg) are considered, spanning material costs from $10 to $104 per kg and densities from 1738 to 12023 kg/m3. We assume a catalyst particle size of 10-6m with 10% coverage. The cost per catalyst surface area is determined by multiplying the cost per kg by the catalyst mass and dividing by the catalyst surface area, yielding $0.02 to $120.2 per m2 of catalyst. Fluxes are computed for the total surface area using Tab. 1 and adjusted for catalyst coverage. Therefore, we estimate the cost per ton of CO2e removed to range from $ to $ for cities and $3.2 to $ for HVAC filters, the latter meeting the $100 per tCO2e target. Further reductions are possible by increasing catalyst coverage or reducing particle size, potentially bringing city-scale applications within target costs.
Discussion
The potential for removing low concentrations of atmospheric pollutants as a mitigation strategy for climate change and public health crises has yet to be fully explored. Surface-based removal technologies have been proposed as a scalable approach to achieving these reductions. For example, it has been shown that DAC may remove CO2 from ambient air at scale [28], catalytic coatings in urban areas can reduce the concentration of volatile organic compounds (VOCs) and NO [43], and HEPA filters can lower PM levels [23]. To evaluate whether wider deployment is warranted, we have developed a framework to assess the potential scale of surface-based atmospheric pollutant removal, leveraging pre-existing airflows in cities, HVAC systems and transportation.
Surface-based pollutant removal technologies in cities, solar farms, HVAC systems and filters could achieve potential atmospheric pollutant removal rates exceeding 1 GtCO2e per year (20–100 year GWP) [20], with optimisation for higher surface fluxes or lower materials costs potentially reducing the minimum cost estimate to under $100 per tCO2e. Building upon the existing capabilities of HVAC systems to remove PM and VOCs [23], by integrating other surface-based pollutant removal technologies, could enhance their cost-effectiveness and scalability. A similar integration of technologies was suggested during the COVID-19 pandemic, when HVAC systems were adapted to mitigate airborne viruses, using technologies such as UV-C light and antimicrobial coatings [48, 51]. This demonstrated the adaptability of HVAC systems and their potential for multifunctional air quality management to address broader environmental challenges. Furthermore, HVAC filters are regularly replaced, potentially making the maintenance of any surface-based removal technologies straightforward and allowing their performance to be sustained over time.
Refining pollutant removal models can not only enhance the accuracy of predictions but also provide insights into optimisation. Adsorption/desorption models, such as Langmuir–Hinshelwood or Eley–Rideal, and multicomponent diffusion models, such as Maxwell–Stefan, can help identify conditions that maximise reaction rates [37, 4]. Similarly, dispersion models, including Gaussian plume, Eulerian and Lagrangian approaches, can enable placement of removal surfaces to enhance the removal efficiency [6]. Lastly, our analyses excluded energy consumption to focus on existing natural and mechanical airflows, assuming thin and smooth coatings with consistent flow resistance. However, rough coatings with increased surface area could enhance turbulence, flow rates to the surface and therefore removal efficiency.
Scaling surface-based atmospheric pollutant removal technologies presents challenges that must be addressed, including material degradation, surface contamination and infrastructure integration. While sorption systems generally maintain consistent pollutant removal efficiencies [46, 42], the catalytic removal of CH4 is variable [50, 32] and may decrease over time or under varying conditions (e.g., relative humidity). Advances in materials and reaction mechanisms can improve these efficiencies and expand the applicability of catalysis. However, large-scale deployment of catalysts may introduce risks, such as the conversion of other species (e.g., VOCs) to more toxic forms.
These findings highlight the potential of integrating surface-based atmospheric pollutant removal technologies into urban and industrial systems to support climate action and protect public health.
Methods
Scaling theory
To characterise atmospheric pollutant transport dynamics across various systems, we define velocity (, m/s), length (, m) and surface area (, m2) scales specific to each flow application (e.g., cities, HVAC systems, transport). We also estimate the number of occurrences () for each flow type to ensure realistic scaling for global estimations. The limitations of this scaling approach and additional context are provided in Supplementary Sec. 3.1.
The thickness of the turbulent () boundary layer (, m) is evaluated using , where the Reynolds number [41] and is the kinematic viscosity (m2/s). The normal pollutant flux (mol/m2/s) at the surface is determined by the diffusive flux, , as the normal velocity at the surface is zero. The normal pollutant flux is simplified here using Fick’s law, which assumes diffusion driven solely by concentration gradients without accounting for molecular interactions or multicomponent effects. This normal pollutant flux depends on the diffusivity (m2/s), concentration (mol/m3) and the boundary layer thickness (m) [13], where the Schmidt number . An effective diffusivity (m2/s) is included to account for enhanced mixing, which alters pollutant transport rates [34].
Using these scales, the streamwise flow rate of pollutants (m3/s) through the boundary layer of each application is given by
The normal flow rate of pollutants (mol/s) to the surface of each application is given by
The formulas in (1)–(2) provide a means to estimate atmospheric pollutant transport through, and to the surface of, global environments. As shown in Supplementary Fig. 1, we calculate the global streamwise flow rate through the boundary layers of the world’s cities (solar farms) using [17, 44, 11].
Empirical relationships
Empirical models can be used to link flow characteristics with the potential for pollutant removal [13]. These relationships have been applied in CH4 and N2O removal studies via photocatalysis [50, 36]. These relationships are evaluated using , and , supplemented with measured values such as wall shear stresses ( in kg/ms2), mass transfer coefficients ( in m/s) and heat transfer coefficients ( in W/m2K). A detailed discussion of these empirical relationships and their applications can be found in Supplementary Sec. 3.2.
Analogies exist between key dimensionless numbers: the Nusselt number (), the Sherwood number () and the drag coefficient (), where is the thermal conductivity (W/mK) and is the fluid density (kg/m3) [13]. These analogies allow for the formulation of the normal pollutant flow rate to the surface of each application as follows
Following Tsopelakou et al. [50], we can use (3) to estimate the streamwise flow rate through each application as
where is the wetted perimeter and is the cross-sectional area of the channel or boundary layer.
Fully-developed velocity and pollutant profiles
We employ established formulas to determine fully-developed velocity and pollutant profiles, including the Monin–Obukhov (M-O) similarity theory for natural airflows and approximate solutions of the Navier-Stokes equations for internal airflows, with explanations provided in Supplementary Sec. 3.3.
For external airflows over cities or solar farms, we utilise M-O similarity theory [41], which incorporates the effects of buoyancy and surface roughness under varying thermal conditions. The normal flow rate of pollutants to the surface of the environment is
where represents the average disturbance height and is the roughness length. For , where is the viscous sublayer thickness, we recover the log-law profile. This log law applies to external airflows over smooth or mildly rough surfaces and can be used to calculate pollutant fluxes near vehicles [34].
For internal airflows within HVAC systems, we approximate solutions of the Navier-Stokes equations and boundary conditions [52]. The normal flow rate of pollutants to the channel walls is given by
where is the channel diameter. For internal airflows through HVAC filters, we adjust the velocity from the duct flow (with velocity , diameter and cross-channel area ) to the pleat flow (with velocity , diameter , filter sheet area , number ). We apply the principle of mass conservation to match average velocities, leading to . The flow rate of pollutants to the pleat channel walls is
The flow rate of pollutants to the fibre sheet is
where is the collection efficiency, related to the permeability and porosity of the filter medium [19].
Energy, power and drag measurements
We now outline formulas to determine the flow rate based on energy, power and drag measurements. More details are given in Supplementary Sec. 3.4.
For external flows over urban areas or solar farms, we establish a control volume with a length and height . In the case of uni-directional, steady-state flow over a surface with a heat flux (W/m2), energy conservation dictates that the energy entering the control volume must balance with the energy exiting it [13]. This relationship can be expressed as
where (m2/s2K) is the specific heat capacity of air at constant pressure and (K) is the temperature difference between the surface and the surrounding atmosphere. The flow rates are related using the scaling theory discussed in Supplementary Sec. 3.1, such that .
In the context of internal flow through HVAC systems, the total power consumed by the fans denoted as (kW) and the specific fan power (SFP) (W/m2/s), can be utilised to evaluate the streamwise flow rate through the duct as follows
where is the leakage factor [40].
For external airflows acting on transportation systems, the streamwise flow rate through the boundary layer can be estimated using the drag force acting on the aeroplane, train or automobile [52]. This is expressed by the equation
where (N) is the drag force.
Industry standards
For internal airflows through HVAC ducting, the streamwise flow rate can be evaluated using three estimates [1]. First, can be expressed in terms of the cubic feet per minute per person (, ft3/min), given by
where (–) is the total number of occupants in the building. Second, can be calculated based on the CFM per unit area (, ft3/min/m2), described by the equation
where (m2) is the total floor area of the building. Finally, can be assessed using the air changes per hour (ACH, s), given by
where (m3) is the volume of air within the building.